Could someone help me with this problem. I am lost. Step by step so I can understand what is done please.
Fast food. If four out of five doctors prefer fast food, then at a convention of 445 doctors, how many prefer fast food?

Answer by rfer(16322)

4 out of 5 = 4/5 = 80%
445*0.80=356